Danny is 40 years old and Jane is 26 years old. How many years ago was Danny thrice as old as Jane?
Answer by Cromlix(4381)

40 - x = 3(26 - x)
40 - x = 78 - 3x
Collect like terms
3x - x = 78 - 40
2x = 38
x = 19
!9 years ago when Danny was 21 years old
and Jane was 7 years old.
Hope this helps.
:-)